The firm is an all-equity firm with assets worth $350 million and 100 million shares outstanding. It plans to borrow $100 million and use these funds to repurchase shares. The firm’s marginal corporate tax is 21%, and it plans to keep its outstanding debt equal to $100 million permanently. If the firm manages to repurchase shares at $4 per share, what is the per share value of equity for the leveraged firm?

$2.71 per share B) $3.5 per share C) $3.61 per share D) $3.71 per share E) $4 per share
Answer #1

According to mm theory value of levered firm will be equal to value of unlevered firm plus pv of tax sheild

Value of assets of levered firm = 350+21%(100) = 371 million

Value of equity of levered equity = assets - debt

= 371-100 = 271

Repurchase price is 4

No of shares repurchased = 100/4 = 25 million

No of shares outstanding = 100 -25 = 75 million

Equity value per share = value of equity/outstanding shares

= 271/75 = 3.61

Option c is correct